### v1

@@ Abstract
This paper examines whether increased awareness can affect racial bias and colorism. For that, we utilize a natural experiment arising from the widespread publicity of [55], which served as an external shock intensifying scrutiny around racial bias in men’s basketball officiating. We investigate refereeing decisions in a similar setting, the Women’s National Basketball Association (WNBA) which is known as a progressive institution with a longstanding commitment to diversity, equity, and inclusion (DEI) policy. We apply state-of-the-art machine learning techniques to systematically predict race and objectively measure player skin tone. Our empirical strategy exploits the quasi-random assignment of referees to games, combined with high-dimensional fixed effects, to estimate the relationship between referee-player racial and skin tone compositions and foul-calling behavior. Our results show no racial bias before the intense media coverage. However, we find evidence of overcorrection, according to which a player earns fewer fouls when facing more referees from the opposite race and skin tone. This overcorrection wore off over time returning to zero-bias levels. We highlight the need to consider baseline levels of bias before applying any prescription with direct relevance to policymakers and organizations given the recent discourse on DEI.
JEL Codes: D12, D91, J15, Z20.
@@ 1  Introduction
Since the seminal work of Gary Becker on the economics of discrimination [11], it has become one of the most researched topics in economics [3]. Discrimination has been found in a variety of settings that include but are not limited to law enforcement [8, 42], justice system [60], housing market [6], labor market [13], and sports [55, 63]. [^5] See [45] for a comprehensive review of empirical literature on the economics of discrimination in the labor market and criminal justice system.
@@
At the time of Becker’s publication in 1957, racial discrimination in the USA was still legal. Today, racial discrimination of any kind is illegal. However, this does not mean that racial discrimination has fully disappeared, and therefore, there is room for increasing awareness of the problem in organizations that have experienced discrimination. Intuitively, such organizations are expected to benefit from increased awareness, which is likely to reduce bias. For example, [55], one of the two studies mostly related to our paper, investigated the racial bias of referees in the National Basketball Association (NBA). The authors found that players earned significantly more fouls from the crew that had more referees of a different skin color. That paper provoked extensive media attention, including front-page coverage in the New York Times and other major media outlets. This widespread coverage was used in [54], the second study that guides our work, as an exogenous shock of increased awareness to investigate whether the bias was reduced after it. The results in [54] showed that this increased awareness indeed eliminated racial bias of the NBA referees. [^6] It is also worth mentioning a recent study by [7] that showed that an increased awareness among Italian teachers reduced the gap between the grades of native versus immigrant students.
@@
However, it is also plausible to assume that there are many (hopefully most) organizations with no existing systematic discrimination. Then the question is how high-profile discriminatory exposures, à-la the above-mentioned NBA case, would affect such organizations. Obviously, the desired outcome would be that an organization with no systematic racial or skin tone discrimination in its organizational culture would not change its behavior during topically relevant and intensified public debates. However, there is the possibility of unintended overcorrection, which refers to the case when individuals or institutions try to avoid or compensate for racial biases. This may lead to an unintentional preference towards a particular racial group. According to [50], such overcorrection may happen because people may be concerned about appearing racially prejudiced.
@@
Even though overcorrection can be closely linked to the widely researched topic of discrimination, the phenomenon of overcorrection has received relatively little attention from economists. To the best of our knowledge, the only economic paper that used natural settings to document overcorrection is [20], who used home advantage in professional cricket as the main driver of bias.777[20] investigated international cricket games where the local umpires were known to be biased in favor of their home nation teams and therefore were replaced by neutral umpires. However, because of the COVID-19 pandemic, the local umpires resumed officiating games of their home nation teams. During these games, the local umpires favored the away teams in a way that is in line with overcorrection. We are not aware of an economic paper on overcorrection in the context of racial bias. In fact, most of the evidence on overcorrection in the racial context comes from experimental psychology. For example, [33] found that when evaluating a poorly written essay, White raters rated the content more positively if they thought that the author was Black than when they believed the author was White. Similarly, [50] found that the majority group members engaged in more positive behavior toward stigmatized or minority group members compared to non-stigmatized or ingroup members. In a more recent study, [38] showed that White patients engaged more with Black providers of medicine than with White or Asian, but there was no difference in engagement between White and Asian providers.
In this paper, we fill a gap in the economic literature by picking a uniquely progressive institution with a longstanding commitment to diversity, equity, and inclusion (DEI), where racial discrimination is very unlikely (which will be explicitly tested), and investigate whether an exogenous increase in awareness of racial discrimination caused overreaction.
@@
More specifically, we build on the efforts of the above-mentioned NBA-related studies of [55] and [54] but expand them in several ways. First, we investigate the possible racial bias in a different, though very similar setting, the Women’s National Basketball Association (WNBA). Since its founding, the WNBA has been synonymous with social justice, actively breaking barriers and advocating for equality [48]. If WNBA’s referees have indeed no systematic racial and skin tone bias (we provide evidence supporting this claim), it would make the WNBA a perfect case to investigate the effect of widespread media coverage on racial discrimination in a non-discriminatory organization.
Second, we use three different measures of race and colorism. Like in [55] and [54], we use human raters to define whether players are Black or not. Moreover, we take advantage of the FairFace algorithm, a state-of-the-art machine learning algorithms trained on a racially balanced dataset of more than 100,000 images that is explicitly designed to mitigate racial biases in facial recognition [40].
Finally, we contribute to the emerging economic literature regarding colorism. For many years, the main approach in the economic literature on discrimination was based on the binary classification of skin color [17]. However, colorism based on skin tone has long been the focus of research within the social sciences [22, 46]. In an increasingly diverse and multiracial world, discrimination based on skin tone has been found to persist even within racial and ethnic groups [51], including African Americans [37, 41], Latin Americans [18, 28], and Asians [47, 70].
Despite this prominence, research on the impacts of skin tone has remained dormant in the economic literature up to recent years. A plausible explanation for the dearth of economics research on colorism could be the lack of objective, quantitative, and scalable approaches to measuring skin tone. Many of the previous papers rely on the National Survey of Black Americans (NSBA) or Multi-City Study of Urban Inequality (MCSUI) [29, 35, 36], which suffer from a lack of representativeness, small sample sizes, imprecise measures of skin tone, and reductive classifications.
Thus, our paper contributes to the growing body of research that employs objective skin tone measurements to uncover discrimination in various settings, including children’s books [4, 62], sports [39], and algorithmic fairness research [56, 67]. For that, we utilize machine learning to measure skin tone continuously and represent it in the CIELAB color space as it offers an objective and standardized method for communicating color information, especially skin color [21, 71]. [^8] Color spaces are mathematical models that include three or four components to represent color information [43]. This system aligns with how humans see color, enabling accurate color measurement and comparison [52]. Thus, we can test not only the effect of the difference in race between the players and the referee crew on the personal foul calls based on a binary definition (Black vs non-Black), but also the effect of the difference in skin tone in a continuous spectrum.
@@
Following the idea presented in [54] and based on close to 13,000 individual fouls, we find that, as expected, the WNBA referees exhibited no racial bias in their foul calls in the years between 2004 and 2006 (prior to the widespread media coverage of the [55] results). However, based on more than 17,000 individual fouls in the years between 2007 and 2010, we find evidence of overcorrection, according to which players earned significantly fewer fouls from the crew that had more referees of a different skin color. In addition, based on the continuous skin tone variable, we were able to detect that referees call fewer fouls the further the distance is between them and the players’ skin tone. In other words, an increased awareness of racial discrimination created an unintended consequence in an organization whose referees were already free of racial bias. Interestingly, in the years between 2011 and 2014, which is another extension to the analyses presented in [54], we find a return to the “pre-awareness” levels of zero bias.
@@ 2  The Women’s National Basketball Association
Like the National Basketball Association (NBA) for men, the WNBA is the premier league for women’s basketball. The similar names underscore the NBA’s direct involvement in creating the WNBA in 1997 as its sister league, initially owning all the teams [69]. Starting in 2002, the NBA sold the franchises, some to their NBA operator counterparts and others to third parties [34]. Despite this interconnectedness, both leagues differ in two relevant ways. First, the players of both leagues have always kept their interests separate through independent player unions that collectively bargain with their respective leagues on issues such as salary and benefits. [^9] The NBA Players Association (https://nbpa.com/) and the WNBA Players Association (https://wnbpa.com/) were formed independently in 1954 and 1998, respectively. Second, referees of the two leagues remained separate until 2017, when the National Basketball Referees Association (NBRA) was brought in to collectively negotiate and represent both leagues’ referees from then on (Women’s National Basketball Association, 2017). [^10] Further notable competition- and gameplay-related differences between the leagues are the number of teams and games, quarter length, ball size, and three-point line distance.
@@ 2.1  Officiating
Each of the approximately 30 referees in the WNBA officiates about 20 regular-season games each year. [^11] Insider knowledge about officiating practices discussed in this section was gained through interviews with long-time WNBA referee Amy Bonner [44] and direct discussions with another veteran referee who wished to remain anonymous. A WNBA season lasts between May and September/October. Referees see each team one to five times per season and rarely work games in the same city without at least a week between assignments. These practical constraints prevent referee assignments from being completely random. However, we follow the procedure presented in Price and Wolfers (2010) to show that for each year in our sample, the number of non-Black referees is unrelated to the number of Black starters (five players who start the game). The results of these tests are presented in Table 1 and the Appendix (Table A.1 for human raters and Table A.2 for skin tone). The results suggest that the assignment decision is unrelated to the racial characteristics of the teams.
@@
  • Notes: Racial classifications determined by FairFace. Human rater classifications are included in the appendix. Each observation is a team×game observation. αFinal column tests: $H_{0}$: # non-Black referees is independent of # Black starters.
@@ 2.2  Social justice
“[Basketball can] help us understand that if we say we care about values like equality, diversity, and inclusion then we must reduce implicit biases not just in the workplace, but at home; not just in our laws, but in our hearts and minds—because that’s what’s expected on the court. And while all of us—athletes and fans alike—face moral choices professionally and personally that can cause discomfort and require risk, today there can no longer be a moral separation between what we do and who we are.” [16]
@@
The quote above emphasizes the WNBA’s commitment to social justice, and its advocacy for equality [48]. The WNBA is unique among professional sports leagues because of its strong emphasis on DEI. [^12] It is evident in who they allow in the league or the treatment of those who do not conform. For example, a former US Senator and team owner, Kelly Loeffler, was forced to sell her stake in the Atlanta Dream following her letter to league commissioner Cathy Engelbert opposing the Black Lives Matter movement [30]. Players even played a decisive role in campaigning for her political opponent, who eventually won [31]. Our interviewed former referee confirms the league’s position that its established commitment to these values obviates the need for bias training.
The WNBA boasts numerous firsts, including the first women to secure a Nike signature shoe deal, the first professional league to launch a pride campaign and dedicate a season to social justice [14], the first professional team to wear “Black Lives Matter” shirts, and the first league to feature married teammates [65].
Moreover, WNBA players consistently speak out on critical issues such as sexuality and equal pay [72]. Mistie Bass, a ten-year veteran, made the point clear, saying, “We are a walking protest at all times as a W.N.B.A. athlete, if you think about it, we have so many different stigmas. We’re just constantly in the fight. I don’t think we have ever not been in a fight for equality, for justice[72]. The players have now institutionalized their fight by forming their own Justice, Equity, Diversity, and Inclusion (JEDI) Committee. Moreover, in solidarity, the league office promotes several initiatives related to civic engagement, voting, and educational resources. [^13] WNBPA Handbook https://wnbpa.com/wp-content/uploads/2024/08/Digital-WNBPA-Union-Handbook.pdf and WNBA Social Justice page https://www.wnba.com/social-justice.
@@ 3.1  Data and measures
We leverage publicly available data spanning 2004 through 2014.141414Box score statistics are from Her Hoop Stats https://herhoopstats.com/. Images are collected from the WNBA League Office https://stats.wnba.com/, gettyimages https://www.gettyimages.com/, or the National Basketball Referee Association https://www.nbra.net/nba-officials/. Relevant biographical information is taken from Basketball Reference https://www.basketball-reference.com/wnba/. As in [55] and [54], we use regular season games for our analyses. For players, our data include detailed game-level statistics, career measures, and individual characteristics such as birthdate, height, weight, and nationality. We also have career statistics and individual characteristics for coaches and referees. Moreover, we systematically collect one front-facing photo per person, prioritizing images from official pre-season media day photoshoots. We conduct a simple Google Images search if no official photo meeting our criteria is available. [^15] This accounts for 18% of our sample size and yields mostly high-quality images. Our key measures, race and skin tone, are derived from the pictures as described below.
Our first approach to measuring race is using the FairFace algorithm [40]. FairFace is a machine learning model trained on a racially balanced dataset of over 108,000 images. As such, and in contrast to other algorithms using face datasets that overrepresent Caucasian faces, it is explicitly designed to mitigate racial biases in facial recognition. [^16] See [40] for algorithm details and performance metrics: https://github.com/dchen236/FairFace. FairFace provides two race classifications, “race4” (White, Black, Asian, Indian) and “race7” (White, Black, Latino_Hispanic, East Asian, Southeast Asian, Indian, Middle Eastern), and assigns probability scores to each category. We used both classifications to create our race variable. An individual is coded as Black (1) if FairFace classifies them as Black in either “race4” or “race7”; otherwise, they are coded as non-Black (0) (see Figure 1).
@@
To validate the FairFace predictions of race, we implemented a second approach closely following [55], i.e., we used human raters to assign race as Black or non-Black, aligning with the US Census definition of Black or African American. [^17] https://www.census.gov/topics/population/race/about.html. One male and one female rater independently rated each player, coach, and referee based on their photos. For each photo, raters indicated “Black” or “non-Black” based on their first impression and then rated their confidence level on a scale of 1 (unsure) to 4 (very confident). We reviewed the ratings and assigned a third rater to resolve discrepancies. We calculate Kappa scores to measure the overall agreement between the FairFace classifications and human raters (see Appendix Table B.1). There was 94.8% agreement, which is “almost perfect agreement” [49].
@@
Our third approach follows the “Image-to-Data Pipeline” in [4], dividing the process into two main parts, i.e., face segmentation and skin tone classification. Face segmentation starts with employing a deep-learning approach called a Fully Connected Convolutional Neural Network (FC-CNN) along with the 2D Face Alignment library to detect 68 facial landmarks. [^18] For more details, see: https://github.com/1adrianb/2D-and-3D-face-alignment?tab=readme-ov-file These landmarks are refined using a convex hull method to generate a precise binary skin mask that isolates the facial region. This process predicts periphery landmarks like the jawline, eyebrows, and lips. From these landmarks, we create a binary skin mask to isolate the facial region before applying this mask to the original image to extract the face and segment the skin for color classification. A Hue Saturation Value (HSV) color space-based filter is applied to ensure that only reliable skin pixels are retained, thus minimizing any lighting artifacts. [^19] The HSV color space represents colors using intuitive values based on tint, saturation, and tone. Importantly, HSV decouples color information (hue and saturation) from brightness (value), making it particularly effective at identifying skin pixels even under uneven lighting conditions [68, 15, 59]. This conversion robustly isolates skin pixels by reducing the influence of shadows and variations in illumination.
@@
Next, we use k-means clustering to identify the 5 most dominant colors of the skin tone pixels and calculate the representative skin tone in the CIELAB (L*a*b*) color space using a weighted average of cluster sizes. In the CIELAB space, L* represents lightness or perceived brightness of a color, while a* and b* represent the red-green and the yellow-blue axes, respectively. All together create a three-dimensional space where each color occupies a unique position based on its L*, a*, and b* coordinates. [^20] Note, this is a departure from the common way to measure skin tone using RGB values [57, 5], because the RGB color space does not reflect human perception when assessing skin color variation [67]. Rather, RGB images combine color and intensity information for each pixel, making it challenging to create a robust method of detecting skin tones due to the impact of lighting and shadows on pixel intensity [19, 59]. Figure 2 displays the distribution of these L*a*b* values across our dataset. For our analysis, we focus on “perceptual skin tint” [4], which is the L* value measured on a scale of 0–100, with 0 representing the darkest and 100 the lightest. Figures 3 and 4 illustrate the pipeline’s predictions on sample images and show the dataset sorted by L*.
@@
Notes: This figure shows clockwise the normal distributions of the L*, a*, and b* values, respectively, obtained as described in [4].
@@
Notes: This figure illustrates the steps of the skin tone analysis for a player (Sue Bird) and a referee (John Butler), divided into two main parts. 1) Face segmentation (left-hand side): The face segmentation process begins by converting headshot images to the red-green-blue (RGB) format, before extracting a face mask using a Fully Connected Convolutional Neural Network (FC-CNN). After this, the FaceAlignment library is used to detect 68 specific facial landmarks, further refined using a Convex Hull method to isolate the targeted facial region. Finally, a Hue Saturation Value (HSV) color space-based filter is applied to ensure that only reliable skin pixels are retained, thus minimizing any lighting artifacts. 2) Skin tone classification (right-hand side): We use k-means clustering to identify the 5 most dominant colors of the skin tone pixels and calculate the representative skin tone in the CIELAB (L*a*b*) color space using a weighted average of cluster sizes. For our further analysis, we focus on “perceptual skin tint” [4], which is the L* value measured on a scale of 0–100, with 0 representing the darkest and 100 the lightest. This process is automated with Python, ensuring reproducibility and efficiency, with further details provided in Appendix Section B.

### v3

@@ Abstract
This paper examines whether increased awareness can affect racial bias and colorism. For this purpose, we utilize a natural experiment arising from the widespread publicity of [price2010], which served as an external shock, intensifying scrutiny of racial bias in men’s basketball officiating. We investigate refereeing decisions in a similar setting, the Women’s National Basketball Association (WNBA), which is known as a progressive institution with a longstanding commitment to diversity, equity, and inclusion (DEI) policy. We apply state-of-the-art artificial intelligence and machine learning techniques to systematically predict race and objectively measure skin tone. Our empirical strategy exploits the quasi-random assignment of referees to games, combined with high-dimensional fixed effects, to estimate the relationship between the racial and skin tone compositions of referees and players, as well as foul-calling behavior. Our results show no significant racial bias before the intense media coverage. However, afterward, we find evidence of overcorrection, according to which a player earns fewer fouls when facing more referees from the opposite race and skin tone. Even though this overcorrection seems to wear off over time, we highlight the need to consider baseline levels of bias before applying any prescription with direct relevance to policymakers and organizations, given the recent discourse on DEI.
JEL Codes: D12, D91, J15, J71, M14, Z20.
@@ 1  Introduction
“A person is aware that they are biased in favor of X; or else they are concerned about the possibility that they might be biased in favor of X; or at least, they are concerned to avoid the appearance of being biased in favor of X. Because of this, they bend over backward in order not to exhibit bias in favor of X. But they overshoot and end up treating X less favorably than they would if they were unbiased.”[kelly2022]
@@
Since the seminal work of Gary Becker on the economics of discrimination [becker1957], it has become one of the most researched topics in economics [abrevaya2012]. Discrimination has been found in a variety of settings that include but are not limited to law enforcement [antonovics2009, knowles2001], justice system [shayo2011], housing market [ahmed2008], labor market [bertrand2004], and sports [price2010, szymanski2000]. [^5] See [lang2020] for a comprehensive review of empirical literature on the economics of discrimination in the labor market and criminal justice system.
@@
At the time of Becker’s publication in 1957, racial discrimination in the USA was still legal. Today, racial discrimination of any kind is illegal. However, this does not mean that racial discrimination has fully disappeared, and therefore, there is room for increasing awareness of the problem in organizations that have experienced discrimination. Intuitively, such organizations are expected to benefit from increased awareness, which is likely to reduce bias. For example, [price2010], one of the two studies mostly related to our paper, investigated the racial bias of referees in the National Basketball Association (NBA). The authors found that players earned significantly more fouls from the crew that had more referees of a different race. That paper provoked extensive media attention, including front-page coverage in the New York Times [schwarz2007] and other major media outlets. The widespread coverage was used in [ppw2018], the second study that guides our work, as an exogenous shock to show that NBA referee racial bias disappeared following increased awareness. [^6] It is also worth mentioning a recent study by [alesina2024] that showed that an increased awareness among Italian teachers reduced the gap between the grades of native versus immigrant students.
@@
However, it is also plausible to assume that many (hopefully most) organizations have few, if any, cases of existing systematic discrimination. Then the question is how high-profile discriminatory exposures, à-la the above-mentioned NBA case, would affect such organizations. Ideally, organizations lacking systematic discrimination would not change their behavior during relevant and intensified public debates. However, unintended overcorrection can occur when individuals or institutions seek to avoid or compensate for past biases, whether perceived or actual. This behavior may lead to an unintentional preference for a particular group. According to [mendes2013], such overcorrection may happen because people may be concerned about appearing racially prejudiced.
@@
Even though overcorrection can be closely linked to the widely researched topic of discrimination, the phenomenon of overcorrection has received relatively little attention from economists. To the best of our knowledge, the only field study to document overcorrection in an empirical setting is [chowdhury2024], who used home advantage in professional cricket as the main driver of bias. They investigated international cricket games where the local umpires were known to be biased in favor of their home nation teams and, therefore, were replaced by neutral umpires. However, because of the COVID-19 pandemic, the local umpires resumed officiating games of their home nation teams. During these games, the local umpires favored the away teams in a way that is in line with overcorrection.
We are not aware of an economic paper on overcorrection in the context of racial bias. Most of the evidence on overcorrection in the racial context comes from lab studies in experimental psychology. For example, [harber1998] found that when evaluating a poorly written essay, White raters rated the content more positively if they thought that the author was Black than when they believed the author was White. Similarly, [mendes2013] found that the majority group members engaged in more positive behavior toward stigmatized or minority group members compared to non-stigmatized or ingroup members. In a more recent study, [howe2022] showed that White patients made more effort to socially engage with Black providers of medicine than with White or Asian. [^7] There was no difference in engagement between White and Asian providers
@@
In this paper, we fill a gap in the literature by investigating whether an exogenous increase in awareness of racial discrimination caused overcorrection within an established institution with a longstanding commitment to diversity, equity, and inclusion (DEI), where explicit discrimination is unlikely. The institutional setting and its policies resemble those adopted later across the public and private sectors following the focus on racial injustice in 2020 [biden2021, wef2022]. More specifically, we build on the efforts of [price2010] and [ppw2018], but expand them in several ways. We investigate possible racial bias in a different, though relevant setting, the Women’s National Basketball Association (WNBA). Since its founding, the WNBA has been synonymous with social justice, actively breaking barriers and advocating for equality [mathewson2020]. If WNBA’s referees indeed have no systematic racial and skin tone bias (we provide evidence supporting this claim), it would make the WNBA a perfect case to investigate the effect of widespread media coverage on racial discrimination in a non-discriminatory organization.
Our paper also contributes to a growing body of research using artificial intelligence and machine learning to uncover discrimination in settings such as children’s books [adukia2023, szasz2023], sports [kamel2023], online job networking [evsyukova2025], and algorithmic fairness research [thong2023]. First, for predictions of race, we use the FairFace algorithm, which is trained on a racially balanced dataset of more than 100,000 images and designed to mitigate racial biases in facial recognition [karkkainenfairface]. Second, we also extend the analysis to skin tone, recognizing that race is subjective [smedley2005] and analyzing only race risks obscuring a key driver of discrimination. Substantial evidence regarding inter- and intra-race skin tone preferences [dixon2017, malema2011, telles2004] confirms that reducing the analysis to race alone oversimplifies the task. Discrimination based on skin tone persists even within racial and ethnic groups [monk2021], including African Americans [hill2002, keith1991], Latin Americans [camposvazquez2021, golashboza2008], and Asians [malema2011, wagatsuma1967]. While the examination of colorism is not necessarily novel [goldsmith2006, hersch2006, hersch2008], existing studies commonly relied on binary skin color classification, presenting a similar binary classification issue as past racial studies [branigan2013]. Moreover, earlier efforts often lacked objective, quantitative, and scalable measures of skin tone measurement. Specifically, previous papers typically relied on surveys such as the National Survey of Black Americans (NSBA) or the Multi-City Study of Urban Inequality (MCSUI), which are limited by factors including lack of representativeness, small sample sizes, and imprecise measurement. We address these critical measurement limitations in past research following the approach proposed by [adukia2023]. More precisely, we measure skin tone continuously on a 0 to 100 scale based on the CIELAB color space, as it offers an objective and standardized method for communicating color information, especially skin color [connolly1997, weatherall1992]. [^8] Color spaces are mathematical models that include three or four components to represent color information [kolkur2017]. This system aligns with how humans see color, enabling accurate color measurement and comparison [moreland2009]. Thus, we can test not only the effect of the difference in race between the players and the referee crew on the personal foul calls based on a binary definition (Black vs non-Black), but also the effect of the difference in skin tone in a continuous spectrum.
@@
As expected, following the methodology presented in [ppw2018] and based on close to 13,000 individual fouls, we do not find significant evidence that WNBA referees exhibited systematic racial bias in their foul calls in the years between 2004 and 2006 (before the widespread media coverage of the [price2010] results). However, based on more than 17,000 individual fouls in the years between 2007 and 2010, we find evidence of overcorrection. When using a binary variable, our results suggest that players receive fewer fouls when officiated by a different race crew. When using the more informative continuous skin tone variable, our estimates substantially gain in precision, showing that referees call fewer fouls the greater the distance is between the referee’s and the players’ skin tone. We interpret this finding with increased awareness of racial discrimination, creating an unintended consequence in an organization whose referees were already free of racial bias. Remarkably, in the years between 2011 and 2014, which is another extension to the analyses also presented in [ppw2018], we find a return to the “pre-awareness” levels.
@@ 2  The Women’s National Basketball Association
Like the National Basketball Association (NBA) for men, the WNBA is the premier league for women’s basketball. The similar names underscore the NBA’s direct involvement in creating the WNBA in 1997 as its sister league, initially owning all the teams [voepel2014]. Starting in 2002, the NBA sold the franchises, some to their NBA operator counterparts and others to third parties [hart2023]. Despite this interconnectedness, both leagues differ in two relevant ways. First, the players of both leagues have always kept their interests separate through independent player unions that collectively bargain with their respective leagues on issues such as salary and benefits. Second, referees of the two leagues remained separate until 2017, when the National Basketball Referees Association (NBRA) was brought in to collectively negotiate and represent both leagues’ referees from then on (Women’s National Basketball Association, 2017). [^9] Other notable competition- and gameplay-related differences between the leagues are the number of teams and games, quarter length, ball size, and three-point line distance.
@@ 2.1  Officiating
Every season, each of approximately 30 referees in the WNBA officiates about 20 regular-season games. [^10] Insider knowledge about officiating practices discussed in this section was gained through interviews with long-time WNBA referee Amy Bonner [lahaye2020] and direct discussions with another veteran referee who wished to remain anonymous. A WNBA season lasts between May and September/October. Referees see each team one to five times per season and rarely work games in the same city without at least a week between assignments. These practical constraints prevent referee assignments from being completely random. However, we follow the procedure presented in Price and Wolfers (2010) to show that for each year in our sample, the number of non-Black referees is unrelated to the number of Black starters (five players who start the game). The results of these tests are presented in Table 1 (as well as Table A.1 for human raters and Table A.2 for skin tone), suggesting that the assignment decision is unrelated to the racial characteristics of the teams.
@@
  • Notes: Testing if the number of non-Black referees is unrelated to the number of black starters for each year in our sample. Racial classifications determined by FairFace. Human rater classifications are included in the appendix. Each observation is a team×game observation. αFinal column tests: $H_{0}$: # non-Black referees is independent of # Black starters.
@@ 2.2  Social justice
“[Basketball can] help us understand that if we say we care about values like equality, diversity, and inclusion then we must reduce implicit biases not just in the workplace, but at home; not just in our laws, but in our hearts and minds—because that’s what’s expected on the court. And while all of us—athletes and fans alike—face moral choices professionally and personally that can cause discomfort and require risk, today there can no longer be a moral separation between what we do and who we are.” [borders2018]
@@
The quote above emphasizes the WNBA’s commitment to social justice, and its advocacy for equality [mathewson2020]. The WNBA is notable among professional sports leagues because of its strong emphasis on DEI. [^11] It is evident in who they accept in the league or the treatment of those who do not conform. For example, a former US Senator and team owner, Kelly Loeffler, was pressured to sell her stake in the Atlanta Dream following her letter to league commissioner Cathy Engelbert opposing the Black Lives Matter movement [graham2021]. Players even played a decisive role in campaigning for her political opponent, who eventually won [gregory2021]. Our interviewed former referee confirms the league’s position that its established commitment to these values obviates the need for bias training.
The WNBA boasts numerous firsts, including the first women to secure a Nike signature shoe deal, the first professional league to launch a pride campaign and dedicate a season to social justice [booker2020], the first professional team to wear “Black Lives Matter” shirts, and the first professional league to feature married teammates [teng2021].
Moreover, WNBA players consistently speak out on critical issues such as sexuality and equal pay [weiner2020]. Mistie Bass, a ten-year veteran, made the point clear, saying, “We are a walking protest at all times as a W.N.B.A. athlete, if you think about it, we have so many different stigmas. We’re just constantly in the fight. I don’t think we have ever not been in a fight for equality, for justice[weiner2020]. The players have now institutionalized their fight by forming their own Justice, Equity, Diversity, and Inclusion (JEDI) Committee. Moreover, in solidarity, the league office promotes several initiatives related to civic engagement, voting, and educational resources. [^12] WNBPA Handbook https://wnbpa.com/wp-content/uploads/2024/08/Digital-WNBPA-Union-Handbook.pdf and WNBA Social Justice page https://www.wnba.com/social-justice.
@@ 3.1  Data and measures
We leverage publicly available data spanning 2004 through 2014.131313Box score statistics are from Her Hoop Stats https://herhoopstats.com/. Images are collected from the WNBA League Office https://stats.wnba.com/, gettyimages https://www.gettyimages.com/, or the National Basketball Referee Association https://www.nbra.net/nba-officials/. Relevant biographical information is taken from Basketball Reference https://www.basketball-reference.com/wnba/. As in [price2010] and [ppw2018], we use regular season games for our analyses. For players, our data includes detailed game-level statistics, career measures, and individual characteristics such as birthdate, height, weight, and nationality. We also have career statistics and individual characteristics for coaches and referees. Moreover, we systematically collect one front-facing photo per person, prioritizing images from official pre-season media day photoshoots. We conduct a simple Google Images search if no official photo meeting our criteria is available. [^14] Images obtained by Google search only account for 18% of our sample size and are mostly of high quality. Our key measures, race and skin tone, are derived from the pictures as described below.
Our first approach to measuring race is using the FairFace algorithm [karkkainenfairface]. FairFace is a machine learning model trained on a racially balanced dataset of over 108,000 images. As such, and in contrast to other algorithms using face datasets that overrepresent Caucasian faces, it is explicitly designed to mitigate racial biases in facial recognition. [^15] See [karkkainenfairface] for algorithm details and performance metrics: https://github.com/dchen236/FairFace. FairFace provides two race classifications, “race4” (White, Black, Asian, Indian) and “race7” (White, Black, Latino Hispanic, East Asian, Southeast Asian, Indian, Middle Eastern), and assigns probability scores to each category. We used both classifications to create our race variable. An individual is coded as Black (1) if FairFace classifies them as Black in either “race4” or “race7”; otherwise, they are coded as non-Black (0) (see Figure 1).
@@
To validate the FairFace predictions of race, we implemented a second approach closely following [price2010], i.e., we used human raters to assign race as Black or non-Black, aligning with the US Census definition of Black or African American. [^16] https://www.census.gov/topics/population/race/about.html. One male and one female rater independently rated each player, coach, and referee based on their photos. For each photo, raters indicated “Black” or “non-Black” based on their first impression and then rated their confidence level on a scale of 1 (unsure) to 4 (very confident). We reviewed the ratings and assigned a third rater to resolve discrepancies. We calculate Kappa scores to measure the overall agreement between the FairFace classifications and human raters (see Appendix Table B.1). There was 94.8% agreement, which is “almost perfect agreement” [mchugh2012].
@@
Our third approach follows the “Image-to-Data Pipeline” in [adukia2023], dividing the process into two main parts, i.e., face segmentation and skin tone classification. Face segmentation starts with employing a deep-learning approach called a Fully Connected Convolutional Neural Network (FC-CNN) along with the 2D Face Alignment library to detect 68 facial landmarks. [^17] For more details, see: https://github.com/1adrianb/2D-and-3D-face-alignment?tab=readme-ov-file These landmarks are refined using a convex hull method to generate a precise binary skin mask that isolates the facial region. This process predicts periphery landmarks like the jawline, eyebrows, and lips. From these landmarks, we create a binary skin mask to isolate the facial region before applying this mask to the original image to extract the face and segment the skin for color classification. A Hue Saturation Value (HSV) color space-based filter is applied to ensure that only reliable skin pixels are retained, thus minimizing any lighting artifacts. [^18] The HSV color space represents colors using intuitive values based on tint, saturation, and tone. Importantly, HSV decouples color information (hue and saturation) from brightness (value), making it particularly effective at identifying skin pixels even under uneven lighting conditions [vezhnevets2003, bora2015, shaik2015]. This conversion robustly isolates skin pixels by reducing the influence of shadows and variations in illumination.
@@
Next, we use k-means clustering to identify the 5 most dominant colors of the skin tone pixels and calculate the representative skin tone in the CIELAB (L*a*b*) color space using a weighted average of cluster sizes. In the CIELAB space, L* represents lightness or perceived brightness of a color, while a* and b* represent the red-green and the yellow-blue axes, respectively. All together create a three-dimensional space where each color occupies a unique position based on its L*, a*, and b* coordinates. [^19] Note, this is a departure from the common way to measure skin tone using RGB values [robst2011, agha2024], because the RGB color space does not reflect human perception when assessing skin color variation [thong2023]. Rather, RGB images combine color and intensity information for each pixel, making it challenging to create a robust method of detecting skin tones due to the impact of lighting and shadows on pixel intensity [chandrappa2011, shaik2015]. Figure 2 displays the distribution of these L*a*b* values across our dataset. For our analysis, we focus on “perceptual skin tint” [adukia2023], which is the L* value measured on a scale of 0–100, with 0 representing the darkest and 100 the lightest. Figures 3 and 4 illustrate the pipeline’s predictions on sample images and show the dataset sorted by L*.
@@
Notes: This figure shows clockwise the normal distributions of the L*, a*, and b* values, respectively, obtained as described in [adukia2023]. It confirms that L* values go, from left to right, dark to light and that a* and b* do not necessarily correspond to skin tint as no obvious light to dark pattern arises.
@@
Notes: This figure illustrates the steps of the skin tone analysis for a player (Sue Bird) and a referee (John Butler), divided into two main parts. 1) Face segmentation (left-hand side): The face segmentation process begins by converting headshot images to the red-green-blue (RGB) format, before extracting a face mask using a Fully Connected Convolutional Neural Network (FC-CNN). After this, the FaceAlignment library is used to detect 68 specific facial landmarks, further refined using a Convex Hull method to isolate the targeted facial region. Finally, a Hue Saturation Value (HSV) color space-based filter is applied to ensure that only reliable skin pixels are retained, thus minimizing any lighting artifacts. 2) Skin tone classification (right-hand side): We use k-means clustering to identify the 5 most dominant colors of the skin tone pixels and calculate the representative skin tone in the CIELAB (L*a*b*) color space using a weighted average of cluster sizes. For our further analysis, we focus on “perceptual skin tint” [adukia2023], which is the L* value measured on a scale of 0–100, with 0 representing the darkest and 100 the lightest. This process is automated with Python, ensuring reproducibility and efficiency, with complete details provided in Appendix Section B.